T1=4C P1=369.01 kpa P2=200kpa find T2 if volumen is contant
Answer #1

Given Data,

P1 = 369.01 kPa

P2 = 200 kPa

T1 = 4° C

V1 = V2 = Constant

Find,

T2 =?

We know that,

P1V1 / T1 =  P2V2 / T2

As Volume is constant, then


P1 / T1 =  P2 / T2

T2 = P2T1 / P1

By putting values,

= 200 x 1000 x 4 / 369.01 x 1000

= 2.1678° C

Corrected up to 4 decimal places.